OpenStudy (anonymous):

Solve the following system of equations by the substitution method. 10x + 10y = 1 x = y - 3 What is the value of y?

OpenStudy (anonymous):

is that the equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

these are also the answer choices -29/20 31/20 -3/2

OpenStudy (anonymous):

10(y-3)+10y=1 10y-30+10y=1 20y-30=1 20y-30+30=1+30 20y=31 y=1.55 see i was asking because you dont get an even anwser..unless let me try to flip it hold on :P

OpenStudy (anonymous):

oh tay c:

OpenStudy (anonymous):

oh its 31/20..lol

OpenStudy (anonymous):

ok that is what i put all my classes are online and math quizes suck :p

OpenStudy (anonymous):

yes they do! i do school online too soo i know how you feel!

OpenStudy (anonymous):

What program do you use?

OpenStudy (anonymous):

flvs

OpenStudy (anonymous):

message me :)